# The environment of TeV halo progenitors

### v1

@@ 2.3 Structure of the surrounding medium
The majority of O and B stars are formed in star clusters (SCs) and OB associations (Wright 2020; Portegies Zwart et al. 2010). A number of massive stars are found in relative isolation in the Galactic field, either because they were created this way, or because they were ejected from clusters due to gravitational interactions, thus becoming runaway stars. In the following, we will handle two different families of progenitors: isolated massive stars and star clusters. The latter category gathers all kinds of groupings irrespective of size or compacity, and it is where most massive stars in our Galaxy are to be found. Zinnecker & Yorke (2007) state that only $2\%$ of B stars and $10-25\%$ of O stars are found isolated in the Galactic field. Motivated by the subsequent analysis of Carretero-Castrillo et al. (2024), we adopted the high-end value of $25\%$ for runaway O stars. Due to the IMF favouring low-mass stars, the impact of the number of runaway O stars is modest, and most runaways are actually B stars. Combining these probabilities we obtain that $10\%$ of massive stars in the mass range we considered are found in isolation while the rest is found in SCs and OB associations.
As a consequence, we considered two different kinds of surrounding medium for the pulsar progenitors: wind-blown bubbles (WBB) and superbubbles (SB), depending on whether the progenitor star is isolated or belongs to a cluster. In both cases, for simplicity, we assumed spherical symmetry for the environment around a star or cluster located at the center. We therefore neglected effects like the proper motions of isolated massive stars or substructures in stellar clusters, both potentially giving rise to asymmetric developments of the environment.
@@ 2.3.1 Wind-blown bubbles around isolated massive stars
The interaction of supersonic stellar winds from isolated massive stars with a typically warm ISM of density $n_{\mathrm{ISM}}$ generates what we call here a wind-blown bubble (WBB). This results in a stratified bubble-like structure (Weaver et al. 1977). In the immediate surroundings of the star, there is the freely expanding wind of speed $u_{\mathrm{w}}$ that ends in a wind termination shock (WTS) at radius $r_{\mathrm{w}}$. After the WTS, one finds a bubble of hot rarefied plasma with density $n_{\mathrm{b}}$ and temperature $T_{\mathrm{b}}$ made of shocked wind and evaporated interstellar material that extends up to radius $r_{\mathrm{b}}$ and occupies most of the volume of the structure. The bubble is bounded by a thin and dense shell of density $n_{\mathrm{shell}}$ found at radius $r_{\mathrm{shell}}$, which consists of all the ISM material that was swept by the bubble during its expansion.
At the end of the MS phase, massive stars enter the RSG phase for $\gtrsim 100\leavevmode\nobreak\ \mathrm{kyr}$ and lose most of their mass in powerful and dense but slow winds. These slow winds and the short duration of the RSG phase are such that they do not impact the extent of the bubble, but they set a new pressure balance in its innermost regions. We therefore assumed that the size of the bubble at the time of the supernova is set from the MS phase only, while the pressure balance is determined from the RSG wind.
The WTS radius $r_{\mathrm{w}}$ is obtained by equating the ram pressure of the wind during the RSG phase and the thermal pressure of the bubble interior at the end of the MS phase:
@@
The RSG winds have a higher ram pressure than MS winds, so the WTS positions we obtained (of order 10 pc) are several times the WTS positions obtained from MS winds (of order 1 pc) as in a number of other similar works (e.g., Dwarkadas 2005). The bubble density $n_{\mathrm{b}}$ and the bubble temperature $T_{\mathrm{b}}$ are taken from Castor et al. (1975) and Weaver et al. (1977), that model the evaporation of the shell matter into the hot bubble:
$n_{\mathrm{b}}=8\times 10^{-3}\leavevmode\nobreak\ {\rm cm}^{-3}$ $\left(\frac{\zeta_{L}}{0.22}\right)^{6/35}\left(\frac{L_{\rm w,MS}}{10^{36}\leavevmode\nobreak\ {\rm erg/s}}\right)^{6/35}$ (14)
×$\times\left(\frac{n_{\rm ISM}}{{\rm cm}^{-3}}\right)^{19/35}\left(\frac{\tau_{\rm w}}{\rm Myr}\right)^{-22/35},$
[…]
×$\times\left(\frac{n_{\rm ISM}}{{\rm cm}^{-3}}\right)^{2/35}\left(\frac{\tau_{\rm w}}{\rm Myr}\right)^{-6/35}.$
The wind time $\tau_{\mathrm{w}}$ that contributes to the expansion of the bubble is defined as $\tau_{\mathrm{w}}=\tau_{\mathrm{MS}}+\tau_{\mathrm{RSG}}$.
We added a corrective factor $\zeta_{L}$ on the luminosity to account for the effect of radiative losses, possibly in relation to mixing and turbulence at the bubble-shell interface. Radiative losses are not included in Weaver et al. (1977) for the evolution of WBBs, leading to bubble sizes bigger than observed (Yadav et al. 2017). Lacking a value specific to WBBs and expecting the behaviour to be similar to SBs, we adopted the value of $\zeta_{L}=22\leavevmode\nobreak\ \%$ proposed in Härer et al. (2023) and based on the observation of SBs.
The bubble radius $r_{\mathrm{b}}$ is given in pc from Weaver et al. (1977):
$r_{\mathrm{b}}=21\leavevmode\nobreak\ \mathrm{pc}\left(\frac{\zeta_{L}}{0.22}\right)^{1/5}\left(\frac{L_{\mathrm{w,MS}}}{10\leavevmode\nobreak\ \mathrm{erg/s}}\right)^{1/5}\left(\frac{\tau_{\mathrm{w,MS}}}{1\leavevmode\nobreak\ \mathrm{Myr}}\right)^{3/5}\left(\frac{n_{\mathrm{ISM}}}{1\leavevmode\nobreak\ \mathrm{cm}^{-3}}\right)^{-1/5}.$ (16)
The wind density profile is obtained by $\rho_{\mathrm{w}}=\dot{M}/4\pi u_{\mathrm{w}}r^{2}$. We recall that the mass $\rho$ and number $n$ densities, that we use interchangeably throughout this paper, are linked by $\rho=\mu m_{\mathrm{p}}n$ with $\mu$ the molecular fraction taken at 1.4 and $m_{\mathrm{p}}$ the proton mass.
After the bubble, the shell contains all the ISM matter that was in the volume of the bubble at the birth of the star and that the expanding bubble swept up. The shell width is assumed to be $\Delta r_{\mathrm{shell}}=1\leavevmode\nobreak\ \mathrm{pc}$ and the volume of the shell is comprised between $r_{\mathrm{b}}$ and $r_{\mathrm{shell}}=r_{\mathrm{b}}+\Delta r_{\mathrm{shell}}$:
@@
The shell and ISM are both considered a warm medium, so from Recchia et al. (2022), we take the corresponding temperatures $T_{\mathrm{shell}}=T_{\mathrm{ISM}}=8000\leavevmode\nobreak\ \mathrm{K}$.
We provide in Table 1 the properties of the WBB at the time of supernova for two stellar progenitors of initial mass $M_{\mathrm{ZAMS}}=8\leavevmode\nobreak\ \mathrm{M}_{\odot}$ and $M_{\mathrm{ZAMS}}=40\leavevmode\nobreak\ \mathrm{M}_{\odot}$.
Progenitor Mass [M] 8 40
MS time [Myr] 37.8 4.6
Ejecta mass [M] 3.1 21.1
RSG mass loss rate [M/yr] $2.9\times 10^{-7}$ $1.6\times 10^{-4}$
RSG wind speed [km/s] 11.0 39.7
MS wind luminosity [erg/s] $1.5\times 10^{31}$ $1.6\times 10^{36}$
RSG Wind radius [pc] 16.4 30.4
Bubble density [cm-3] $5.3\times 10^{-5}$ $1.4\times 10^{-3}$
Bubble temperature [K] $6.2\times 10^{4}$ $1.3\times 10^{6}$
Bubble radius [pc] 20.1 57.6
Shell density [cm-3] 4.5 13.5
Shell radius [pc] 21.1 58.6
Shell Mass [M] 841 24000
@@
2.3.2 Superbubbles around massive star clusters
SBs can be described in the same theoretical framework as WBBs (Weaver et al. 1977; Härer et al. 2023), replacing the wind properties of a single star by the collective mechanical power of all stars of the cluster. As a result, SBs extend over hundreds of parsecs instead of tens of parsecs for WBBs.
In this work, we neglected any structure in the stellar cluster or its spatial extent and assumed that SBs are the result of a mechanical outflow emanating from a central point. For each SB, we first sampled the total cluster mass $M_{\mathrm{cl}}$ between $10^{3}\leavevmode\nobreak\ \mathrm{M}_{\odot}\lesssim M_{\mathrm{cl}}\lesssim 10^{5}\leavevmode\nobreak\ \mathrm{M}_{\odot}$ from the following young massive star cluster IMF (Portegies Zwart et al. 2010):
@@ 2.3.2 Superbubbles around massive star clusters
with the cutoff mass $M_{\mathrm{cl,cut}}=2\times 10^{5}\leavevmode\nobreak\ \mathrm{M}_{\odot}$ for spiral galaxies. We then populated the SC with stars of mass $M_{\mathrm{ZAMS}}$ taken from a Salpeter IMF with index $\alpha=2.35$:
$\frac{\mathrm{d}N}{\mathrm{d}M_{\mathrm{ZAMS}}}\propto M_{\mathrm{ZAMS}}^{-\alpha}\exp(-M_{\mathrm{\mathrm{ZAMS},cut}}/M_{\mathrm{ZAMS}})$ (19)
for masses between $0.1\leavevmode\nobreak\ \mathrm{M}_{\odot}<M_{\mathrm{ZAMS}}<150\leavevmode\nobreak\ \mathrm{M}_{\odot}$ and adopting a cutoff mass of $M_{\mathrm{\mathrm{ZAMS},cut}}=0.35\leavevmode\nobreak\ \mathrm{M}_{\odot}$ (Larson 1998). We sampled from this distribution until the cluster’s mass is reached.
The SC properties needed for modelling the SB are its mechanical luminosity $L_{\mathrm{w,SC}}$, and its collective wind mass loss rate $\dot{M}_{\mathrm{SC}}$ and speed $u_{\mathrm{w,SC}}$. The first two quantities are simple sums of the luminosity $L_{\mathrm{w}}$ and mass loss rate $\dot{M}$ of individual massive stars $i$:
@@
while the SC collective wind speed $u_{\mathrm{w,SC}}$ is linked to the wind speeds $u_{\mathrm{w}}$ and mass loss rates of individual stars using momentum conservation as in Menchiari et al. (2024):
@@
As stellar wind properties of individual stars, we adopted a different approach than that followed in Sect. 2.3.1 for the WBBs, for the sake of simplicity (because the superposition of different winds from a high number of stars at different evolutionary stages is a complex problem). First, we considered O and B stars with masses $>2\leavevmode\nobreak\ \mathrm{M}_{\odot}$ (and not just for $8\leavevmode\nobreak\ \mathrm{M}_{\odot}<M_{\mathrm{ZAMS}}<40\leavevmode\nobreak\ \mathrm{M}_{\odot}$, which was motivated by the ability to form a pulsar), as this broader mass range yields a better account of the total energetics of the cluster. Second, we restricted the wind properties to those of the MS stage, which we picked from Seo et al. (2018). Last, we assumed that the collective mechanical power of the SC remains constant and equal to its initial value, when all its member stars are in the MS stage. Thus, we did not account for any change in luminosity when stars leave the MS and enter RSG or WR stages, and we also missed the decrease in luminosity when the massive stars have exploded. Conversely, we did not consider the energy input from supernova explosions. A more realistic description of a SB including these effects would have to rely on hydrodynamical simulations of the time-dependent outflows from an ageing star population, but this complex modelling goes beyond the scope of our statistical work.
Finally, since SCs are born in giant molecular clouds (GMCs) with typical gas number density $n_{\mathrm{GMC}}\sim 100\leavevmode\nobreak\ \mathrm{cm}^{-3}$ (Parizot et al. 2004), we took this value for the density of the ambient medium surrounding the SB, thus setting $n_{\mathrm{ISM}}=n_{\mathrm{GMC}}$.
Out of all the stars able to produce pulsars (we recall $8\leavevmode\nobreak\ \mathrm{M}_{\odot}<M_{\mathrm{ZAMS}}<40\leavevmode\nobreak\ \mathrm{M}_{\odot}$) in the cluster, we randomly picked one and computes its MS time. Contrary to the WBB case, when the selected massive star that produces a pulsar dies, the rest of the cluster can still act as an engine for the SB so we evolved the system after the SN of the selected star until the pulsar reaches the shell and escapes.

### v2

@@ 2.3 Structure of the surrounding medium
The majority of O and B stars are formed in star clusters (SCs) and OB associations (Wright 2020; Portegies Zwart et al. 2010). A number of massive stars are found in relative isolation in the Galactic field, either because they were created this way, or because they were ejected from clusters due to gravitational interactions, thus becoming runaway stars. In the following, we handle two different families of progenitors: isolated massive stars and SCs. The latter category gathers all types of groupings irrespective of size or compactness, and it is where most massive stars in our Galaxy are to be found. Zinnecker and Yorke (2007) state that only $2\%$ of B stars and $10-25\%$ of O stars are found isolated in the Galactic field. Motivated by the subsequent analysis of Carretero-Castrillo et al. (2023), we adopted the high-end value of $25\%$ for runaway O stars. Due to the IMF favouring low-mass stars, the impact of the number of runaway O stars is modest, and most runaways are actually B stars. Combining these probabilities we find that $10\%$ of the massive stars in the mass range we considered are found in isolation, while the rest is found in SCs and OB associations.
As a consequence, we considered two different types of surrounding medium for the pulsar progenitors: wind-blown bubbles (WBB) and superbubbles (SB), depending on whether the progenitor star is isolated or belongs to a cluster. In both cases, for simplicity, we assumed spherical symmetry for the environment around a star or cluster located at the centre. We therefore neglected effects such as the proper motions of isolated massive stars or the substructures in stellar clusters, both potentially giving rise to asymmetric developments of the environment.
@@ 2.3.1 Wind-blown bubbles around isolated massive stars
The interaction of supersonic stellar winds from isolated massive stars with a typically warm ISM of density $n_{\mathrm{ISM}}$ generates what we call here a WBB. This results in a stratified bubble-like structure (Weaver et al. 1977). In the immediate surroundings of the star, freely expanding wind of speed $u_{\mathrm{w}}$ ends in a wind termination shock (WTS) at radius $r_{\mathrm{w}}$. After the WTS, one finds a bubble of hot rarefied plasma with density $n_{\mathrm{b}}$ and temperature $T_{\mathrm{b}}$ made of shocked wind and evaporated interstellar material that extends up to radius $r_{\mathrm{b}}$ and occupies most of the volume of the structure. The bubble is bounded by a thin and dense shell of density $n_{\mathrm{shell}}$ found at radius $r_{\mathrm{shell}}$, which consists of all the ISM material that was swept by the bubble during its expansion.
At the end of the MS phase, massive stars enter the RSG phase for $\gtrsim 100\penalty 10000\ \mathrm{kyr}$ and lose most of their mass in powerful and dense but slow winds. These slow winds and the short duration of the RSG phase are such that they do not impact the extent of the bubble, but they set a new pressure balance in its innermost regions. We therefore assumed that the size of the bubble at the time of the supernova is set from the MS phase only, while the pressure balance is determined from the RSG wind.
The WTS radius $r_{\mathrm{w}}$ was obtained by equating the ram pressure of the wind during the RSG phase and the thermal pressure of the bubble interior at the end of the MS phase:
@@
In general, a star of a certain mass in our range exhibits more powerful winds in the RSG than the MS phase, so the WTS positions we obtained (of the order of 10 pc) are several times the WTS positions obtained from MS winds (of order 1 pc) as in a number of other similar studies (e.g. Dwarkadas 2005). The bubble density $n_{\mathrm{b}}$ and the bubble temperature $T_{\mathrm{b}}$ were taken from Castor et al. (1975) and Weaver et al. (1977), who model the evaporation of the shell matter into the hot bubble:
$n_{\mathrm{b}}=8\times 10^{-3}\penalty 10000\ {\rm cm}^{-3}$ $\left(\frac{\zeta_{\mathrm{L}}}{0.22}\right)^{6/35}\left(\frac{L_{\rm w,MS}}{10^{36}\penalty 10000\ {\rm erg/s}}\right)^{6/35}$ (15)
$\times\left(\frac{n_{\rm ISM}}{{\rm cm}^{-3}}\right)^{19/35}\left(\frac{\tau_{\rm w}}{\rm Myr}\right)^{-22/35},$
$T_{\mathrm{b}}=\ 1.46\times 10^{6}\penalty 10000\ {\rm K}$ $\left(\frac{\zeta_{\mathrm{L}}}{0.22}\right)^{8/35}\left(\frac{L_{\rm w,MS}}{10^{36}\penalty 10000\ {\rm erg/s}}\right)^{8/35}$ (16)
$\times\left(\frac{n_{\rm ISM}}{{\rm cm}^{-3}}\right)^{2/35}\left(\frac{\tau_{\rm w}}{\rm Myr}\right)^{-6/35}.$
The wind time $\tau_{\mathrm{w}}$ contributing to the expansion of the bubble is defined as $\tau_{\mathrm{w}}=\tau_{\mathrm{MS}}+\tau_{\mathrm{RSG}}$.
We added a corrective factor $\zeta_{\mathrm{L}}$ on the luminosity to account for the effect of radiative losses, possibly in relation to mixing and turbulence at the bubble-shell interface. Radiative losses are not included in Weaver et al. (1977) for the evolution of WBBs, leading to bubble sizes bigger than observed (Yadav et al. 2017). Lacking a value specific to WBBs and expecting the behaviour to be similar to SBs, we adopted the value of $\zeta_{\mathrm{L}}=22\penalty 10000\ \%$ proposed in Härer et al. (2023) and based on the observation of SBs.
The bubble radius $r_{\mathrm{b}}$ is given in parsecs from Weaver et al. (1977):
$r_{\mathrm{b}}=21\penalty 10000\ \mathrm{pc}\left(\frac{\zeta_{\mathrm{L}}}{0.22}\right)^{1/5}\left(\frac{L_{\mathrm{w,MS}}}{10^{36}\penalty 10000\ \mathrm{erg/s}}\right)^{1/5}\left(\frac{\tau_{\mathrm{w,MS}}}{1\penalty 10000\ \mathrm{Myr}}\right)^{3/5}\left(\frac{n_{\mathrm{ISM}}}{1\penalty 10000\ \mathrm{cm}^{-3}}\right)^{-1/5}.$ (17)
The wind density profile is obtained by $\rho_{\mathrm{w}}=\dot{M}/4\pi u_{\mathrm{w}}r^{2}$. We recall that the mass $\rho$ and number $n$ densities, used interchangeably throughout this paper, are linked by $\rho=\mu m_{\mathrm{p}}n$ with $\mu$ the molecular fraction taken at 1.4 and $m_{\mathrm{p}}$ the proton mass.
After the bubble, the shell contains all the ISM matter that was in the volume of the bubble at the star’s birth and that was swept up by the expanding bubble. The shell width is assumed to be $\Delta r_{\mathrm{shell}}=1\penalty 10000\ \mathrm{pc}$, and the volume of the shell lies between $r_{\mathrm{b}}$ and $r_{\mathrm{shell}}=r_{\mathrm{b}}+\Delta r_{\mathrm{shell}}$:
@@
The shell and ISM are both considered a warm medium, so from Recchia et al. (2022), we take the corresponding temperatures $T_{\mathrm{shell}}=T_{\mathrm{ISM}}=8000\penalty 10000\ \mathrm{K}$. We provide in Table 1 the properties of the WBB at the time of supernova for two stellar progenitors of initial mass $M_{\mathrm{ZAMS}}=8\penalty 10000\ \mathrm{M}_{\odot}$ and $M_{\mathrm{ZAMS}}=20\penalty 10000\ \mathrm{M}_{\odot}$.
Progenitor Mass [M] 8 20
MS time [Myr] 37.8 9.6
Ejecta mass [M] 3.1 10.4
RSG mass loss rate [M/yr] $2.9\times 10^{-7}$ $1.2\times 10^{-5}$
RSG wind speed [km/s] 11.0 23.4
MS wind luminosity [erg/s] $1.5\times 10^{31}$ $5.1\times 10^{34}$
RSG Wind radius [pc] 16.4 17.2
Bubble density [cm-3] $5.3\times 10^{-5}$ $5.0\times 10^{-4}$
Bubble temperature [K] $6.2\times 10^{4}$ $5.0\times 10^{5}$
Bubble radius [pc] 20.1 45.4
Shell density [cm-3] 4.5 10.6
Shell radius [pc] 21.1 46.4
Shell Mass [M] 841 12793
@@
2.3.2 Superbubbles around massive SCs
Superbubbles can be described in the same theoretical framework as WBBs (Weaver et al. 1977; Härer et al. 2023), replacing the wind properties of a single star by the collective mechanical power of all stars of the cluster. As a result, SBs extend over hundreds of parsecs instead of tens of parsecs for WBBs.
In this work, we neglected any structure in the stellar cluster or its spatial extent and assumed that SBs are the result of a mechanical outflow emanating from a central point. For each SB, we first sampled the total cluster mass $M_{\mathrm{cl}}$ between $10^{3}\penalty 10000\ \mathrm{M}_{\odot}\lesssim M_{\mathrm{cl}}\lesssim 10^{5}\penalty 10000\ \mathrm{M}_{\odot}$ from the following young massive SC IMF (Portegies Zwart et al. 2010),
@@ 2.3.2 Superbubbles around massive SCs
with the cutoff mass $M_{\mathrm{cl,cut}}=2\times 10^{5}\penalty 10000\ \mathrm{M}_{\odot}$ for spiral galaxies. We then populated the SC with stars of mass $M_{\mathrm{ZAMS}}$ taken from a Salpeter IMF with index $\alpha=2.35$,
$\frac{\mathrm{d}N}{\mathrm{d}M_{\mathrm{ZAMS}}}\propto M_{\mathrm{ZAMS}}^{-\alpha}\exp(-M_{\mathrm{\mathrm{ZAMS},cut}}/M_{\mathrm{ZAMS}}),$ (20)
for masses between $0.1\penalty 10000\ \mathrm{M}_{\odot}<M_{\mathrm{ZAMS}}<150\penalty 10000\ \mathrm{M}_{\odot}$, and adopted a cutoff mass of $M_{\mathrm{\mathrm{ZAMS},cut}}=0.35\penalty 10000\ \mathrm{M}_{\odot}$ (Larson 1998). We sampled from this distribution until the cluster’s mass was reached.
The SC properties needed for modelling the SB are its mechanical luminosity $L_{\mathrm{w,SC}}$, its collective wind mass loss rate $\dot{M}_{\mathrm{SC}}$, and speed $u_{\mathrm{w,SC}}$. The first two quantities are simple sums of the luminosity $L_{\mathrm{w}}$ and mass loss rate $\dot{M}$ of individual massive stars $i$:
@@
In contrast, the SC collective wind speed $u_{\mathrm{w,SC}}$ is linked to the wind speeds $u_{\mathrm{w}}$ and mass loss rates of individual stars using momentum conservation as in Menchiari et al. (2024):
@@
In 1D, all the stellar winds contribute fully to the wind luminosity, while in reality some energy is lost through wind interactions. The losses remain marginal when considering the bubble size. This was shown in Vieu et al. (2024) where the bubble still evolves by following the Weaver et al. (1977) model, even if they show an absence of WTS around the OB association. As stellar wind properties of individual stars, we adopted a different approach than that followed in Sect. 2.3.1 for the WBBs, for the sake of simplicity (because the superposition of different winds from a high number of stars at different evolutionary stages is a complex problem). First, we considered O and B stars with masses $2\penalty 10000\ \mathrm{M}_{\odot}$ (and not just for $8\penalty 10000\ \mathrm{M}_{\odot}<M_{\mathrm{ZAMS}}<20\penalty 10000\ \mathrm{M}_{\odot}$, which was motivated by the ability to form a pulsar), as this broader mass range yields a better account of the total energetics of the cluster. Second, we restricted the wind properties to those of the MS stage, which we picked from Seo et al. (2018). Last, we assumed that the collective mechanical power of the SC remains constant and equal to its initial value, when all its member stars are in the MS stage. Thus, we did not account for any change in luminosity when stars of the cluster leave the MS and enter RSG or WR stages, and we also missed the decrease in luminosity when the massive stars exploded. Conversely, we did not consider the energy input from SN explosions. Nevertheless, this contribution was assessed in Vieu et al. (2022), and the authors showed that the mechanical power of a cluster is roughly constant over the first tens of millions of years, which suggests that our assumption of a constant luminosity is reasonable. A more realistic description of a SB including these effects would have to rely on hydrodynamical simulations of the time-dependent outflows from an ageing star population, but this complex modelling goes beyond the scope of our statistical work.
Finally, Parizot et al. (2004) found that SCs are born in giant molecular clouds (GMCs) with a typical gas number density $n_{\mathrm{GMC}}\sim 100\penalty 10000\ \mathrm{cm}^{-3}$. We took this value for the density of the ambient medium surrounding the SB, thus setting $n_{\mathrm{ISM}}=n_{\mathrm{GMC}}$.
Out of all the stars able to produce pulsars (we recall $8\penalty 10000\ \mathrm{M}_{\odot}<M_{\mathrm{ZAMS}}<20\penalty 10000\ \mathrm{M}_{\odot}$) in the cluster, we randomly picked one and computed its MS time. Contrary to the WBB case, when the selected massive star that produces a pulsar dies, the rest of the cluster can still act as an engine for the SB; therefore, we evolved the system after the SN of the selected star until the pulsar reached the shell and escaped.